OpenStudy (anonymous):

How to solve -(x-10)=7

OpenStudy (anonymous):

-x+10=7 -x=-3 x=3

OpenStudy (anonymous):

Multiply parentheses by -1 then subtract 10 from both sides of the equal sign then make -x positive.